OpenStudy (anonymous):

the medium side of a right triangle is 7 more than the shortest side. The longest side is 7 less than 3 times the shortest side. Find the length of the shortest side of the triangle

OpenStudy (jdoe0001):

 \(\bf c^2=a^2+b^2\implies (3x-7)^2=(x+7)^2+(x)^2\) solve for "x"